Solve the problem. The profit that the vendor makes per day by selling x pretzles is given by the function P(x) = -0.002x squared + 1.2x - 300. Find the number of pretzels that must be sold to maximize profit.
Answer by stanbon(75887)

Solve the problem. The profit that the vendor makes per day by selling x pretzles is given by the function P(x) = -0.002x squared + 1.2x - 300. Find the number of pretzels that must be sold to maximize profit.
-----------------
Max occurs when x = -b/2a
= -1.2/(2*-0.002) = 300 (# of pretzels that must be sold)
---------
Cheers,
Stan H.